[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Mon Jan 28 08:35:02 EST 2019

Logentry Text:
Over the weekend, all cavities in 1L20 and most cavities in 1L07 were helium processed.  It is also possible that cavities 1L06 and 1L21 may also need processing because of helium bleed through.  SRF plans to initiate the 30K cycle sometime tomorrow afternoon with the hope that RF will be restored as early as Wednesday morning.  It is difficult to say how much progress has been made so far in terms of achieving the desired gradient of 2.1 GeV/pass.  We will know more as work continues early on this week.

There will be no RC meeting today.

Tasks to be completed in the Hall prior to lock up:

Aerogel installation (Vlad, complete, https://logbooks.jlab.org/entry/3645382, https://logbooks.jlab.org/entry/3645762)
HGC installation (Brad, complete, https://logbooks.jlab.org/entry/3645628)
Moller magnet cool down test (Dave G., complete, https://logbooks.jlab.org/entry/3645289)
Installation of lead shielding for J/Psi radiator (Techs, complete)
Beam line and scattering chamber pump down (Dave M., complete, https://logbooks.jlab.org/entry/3645404)
Target cool down with helium gas (Dave M., complete)
Current load test of HMS & SHMS magnets (Holly, complete, https://logbooks.jlab.org/entry/3645602)
Survey & alignment of downstream big beam line, harps, and target (S&A group, ?) 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3645961
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190128/6ce0bd3f/attachment.html>

More information about the Hallc_running mailing list